OpenStudy (anonymous):

The formula for the area of a trapezoid is A=1/2h(a+b), where h represents the height and a and b represent the length of the bases. Solve for h.

OpenStudy (anonymous):

Okay, you just need to move everything on the right side of the equation (except h) to the left side of the equation.

jhonyy9 (jhonyy9):

so this you can rewriting like,in the form of : h(a+b) A = -------- 2 multiplie both sides by 2 and will get 2A = 2h(a+b)/2 simplifie in the right side by 2 and will get 2A=h(a+b) so now divide both sides by (a+b) 2A/(a+b) = h(a+b)/(a+b) simplifie in the right side by (a+b) and will get 2A/(a+b) =h so this is the answer hope do you understand it sure right good luck bye